Sunday 28th February

Cryo intervention from midnigth to Monday noon
                           Access from 7am to noon possible on Monday

Overnight Saturday - Sunday: injection setup; beam threading

Day: splashes for experiments; BI checks; establishing multi turns and capture beams with RF

8:30 meeting slides

Saturday 27th February

Overnight Friday-Saturday problems related to free wheel diodes of triplets after the triplet circuits tripped during pre-cycle.

Filled in QT12 QPS intervention and water leak intervention in IR1 in parerallel to diode intervention. All interventions done by 3pm.

Problem related to current lead heater for Q6L2B2; no expert available until 7pm; problem fixed by 9pm.

Preparing machine for injection settings around 10pm.

QPS trips S67 at 10pm after one hour at injection setting.

Trying to recover by midnight.

Friday 26th February

Overnight Thursday - Friday powering tests (pre-cycles and ramp)

Access required this morning for beam interlock problem in point 6 (after patrol - glitch on access door overnight). Aim to closed up for midday, interlock tests followed by beam dump energy tracking tests. Full machine precycle before attempting to inject beam - late afternoon.

Thursday 25th February

21:00 Access and various during the day. Closed up early evening, full machine set to 450 GeV. Beam interlock loops closed for beam 1, minor issue for beam 2. Initial beam dump energy tracking tests performed.

09:00 Following yesterday 50 magnet quench incident which reveal a feature in the Quench Protection logic, the main bend current has been limited to 2 kA until a full analysis has been performed and compensatory measures implemented. (The quenching is not so dangerous in itself but it does unecessarily stress the system).

We will carry on the beam commissioning program as planned with the above limitation in place for the moment. Still hope to put beam in tomorrow.

Wednesday 24th February

23:00: nQPS triggered a 50 magnet quench in sector 78 following non-standard sequence of events. Situation to fully understood before proceding - all commissioning of main dipoles on hold overnight.

18:00: Hope to finish off S78 and S81 this evening. Powering tests overnight. Access needed tomorrow plus beam dump tests and full OP checks through evening and overnight. Aim to inject around 8 a.m. Friday morning if things go well.

Tuesday 23rd February

Sectors 12, 23, 34, 45, 56 and 67 more-or-less good to go.
S78 - 5 kA pyramids to finish
S81 - 2 kA & 5 kA pyramids to finish (following interventions in the tunnel today).

Estimate finishing S81 tomorrow afternoon if things go well. This needs to be followed by final powering tests etc. Thursday daytime for fullscale operations' tests and beam dump energy tracking qualification. Beam Thursday evening at the earliest.

Monday 22nd February

Sector 12, 23, 34, 45, 56 in good shape - still some minor issues.
S67 - interventions today, heading for 6 kA
S78 - interventions today, 2 kA
S81 - nQPS tests progressing well

Beam Interlock System connected today, as was the beam dumps to the access system, and beam energy tracking system to the main bends.

Sunday 21st February

Sectors 12, 23 , 34 - heat runs (last step)
S45 final tests started - minor issues
S56 final tests
S67 ongoing - 2 kA
S78 should finish 2 kA tests today
S81 Good progress, up to 2 kA

Saturday 20th February

Sectors 12, 23, 34, 45, 56 - final powering tests - minor issues.
S67 - 2 kA - intervention required
S78 - ramped to injection - nQPS commissioning
S81 - nQPS commissioning

Friday 19th February

Transfer line tests overnight successful. Full program of measurements performed - things look good.

Sectors 12, 23, 34 - final powering tests - some minor issues
S45 - 5 kA pyramid done - analysis pending
S56 - final powering tests advancing well
S67 - 2 kA - issues. UPS problem - no powering until later this afternoon
S78 - nQPS tests ongoing.
S81 - cryogenics not yet back after transformer intervention (2 kA tests to be continued)

Thursday 18th February

Preparation overview [ppt]

Sectors 12, 23, 34, PGC progressing (although issue with quenching magnets during fast power abort tests to be followed-up).
S45 - 4 kA
S56 - start PGC
S67 - UPS problem to be fixed. 2 kA.
S78 & 81: cryogenics recovering from transformer intervention - back later today. 2 kA tests.

TI2 and TI8 tests tonight - for program see above presentation.

Wednesday 17th February

Start of transformer repair in point 8. DSO (Departmental Safety Officer) tests in TI2 - no powering in sectors 12, 23, 34. No access to point 2.

Sectors 12, 23, 34 in final stages of Powering Group of Circuit (PGC) tests. Some debugging.

Sector 45, 56, 67, 78, 81 all powered - nQPS commissioning and general debugging ongoing.

Monday 15th February

Hardware commissioning continues with sectors 12, 23, 34 ungoing their final powering tests. The other 5 sectors are still in the process of new Quench Protection system commissioning.

Plans for the week:

No beam in the LHC before next week.